OpenStudy (anonymous):

Change to trigonometric form; then find the indicated power. Leave your result in trigonometric form: r(cos theta+i sin theta). -1 - i in trigonometric form = a. 2(cos225+isin225) b. 2(cos135+isin135) c. sqrt2(cos225+isin225) d. sqrt2(cos135+isin135)

OpenStudy (dumbcow):

use these formulas for any complex number : a+bi \[r = \sqrt{a^{2}+b^{2}}\] \[\tan \theta = \frac{b}{a}\]

OpenStudy (dumbcow):

also you need to identify which quadrant you are in to get the right angle imagine you are plotting the point (a,b)

OpenStudy (dumbcow):

hope i didn't lose you...do you understand the formula and what (a,b) refer to for this problem

OpenStudy (anonymous):

Oh, sorry, bad connection :/ But yea, so r=1.414?

hartnn (hartnn):

yes,right...which is sqrt 2

OpenStudy (anonymous):

So now... ?

hartnn (hartnn):

now u find tan theta=b/a..and hence find theta

OpenStudy (anonymous):

I think I got it. C.?

hartnn (hartnn):

yes C it is :)

OpenStudy (anonymous):

Thank you!